### v2

Abstract
Existing language models (LMs) often exhibit a Western-centric bias and struggle to represent diverse cultural knowledge. Previous attempts to address this rely on synthetic data and express cultural knowledge only in English. In this work, we study whether a small amount of human-written, multilingual cultural preference data can improve LMs across various model families and sizes. We first introduce CARE, a multilingual resource of 24.1k responses with human preferences on 2,580 questions about Chinese and Arab cultures, all carefully annotated by native speakers and offering more balanced coverage. Using CARE, we demonstrate that cultural alignment improves existing LMs beyond generic resources without compromising general capabilities. Moreover, we evaluate the cultural awareness of LMs, native speakers, and retrieved web content when queried in different languages. Our experiment reveals regional disparities among LMs, which may also be reflected in the documentation gap: native speakers often take everyday cultural commonsense and social norms for granted, while non-natives are more likely to seek out and document them actively. CARE is publicly available at https://github.com/Guochry/CARE (we plan to add Japanese data in the near future).
CARE: Aligning Language Models for Regional Cultural Awareness
[…]
1 Introduction
The impressive capabilities of multilingual language models (LMs) lead to their widespread use across various cultural groups Ouyang et al. (2022); Dubey et al. (2024). However, many studies have shown that LMs often lack knowledge of diverse cultures and are biased towards Western culture Naous et al. (2023); AlKhamissi et al. (2024); Naous and Xu (2025). While much focus has been on demonstrating these cultural biases, less effort Li et al. (2024a, b); Kirk et al. (2024) has been made to address them. Existing works have largely been limited to expressing cultural knowledge in only English or generating synthetic data with existing LMs for fine-tuning.
Figure 1: Example LM responses to culture-specific questions in Chinese and Arabic. The base LM (Refer to caption) Llama3.1-8B-Instruct fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption) and Arab (Refer to caption) cultures.
In this work, we explore the benefits of learning from multilingual multicultural human preference data, and demonstrate how even a modest dataset of 2.6k high-quality human-written samples can significantly enhance LM’s cultural awareness across various model families and sizes.
Specifically, we first introduce the Cultural Alignment REsource (CARE [Uncaptioned image]), comprising 24.1k responses (one human-written and nine LM-generated per question) with native human preference using a 10-point scale on 2580 multilingual questions focusing on Chinese and Arab cultures (825 and 697 in respective languages; additional 1,058 covering other cultures). Compared with existing datasets Yin et al. (2022); Palta and Rudinger (2023); Shi et al. (2024); Chiu et al. (2024b), our data offers more balanced coverage and authentic viewpoints, especially for topics such as social norms and cultural commonsense, where previous works have limited data coverage. To achieve this, we have native speakers both re-annotate existing data resources, and curate new data on those rarely documented yet deeply embedded aspects of daily life.
Using CARE, we demonstrate that multilingual human cultural preference enhances LMs’ cultural awareness through alignment strategies, such as direct preference optimization (DPO, Rafailov et al., 2024), without compromising general capabilities. This cultural alignment proves beneficial beyond standard human preference datasets, with consistent gains across both chain-of-thought (CoT) and role-play prompting settings. We further analyze post-alignment performance across different LMs and find that for cultural alignment to be effective, base LMs need a foundational level of cultural knowledge. That is, cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. We then examine the impact of cross-cultural alignment, and notice that aligning models with data from multiple cultures results in greater improvements than using single-culture data. This highlights the benefits of expanding geographical coverage in LM alignment.
Moreover, we evaluate the cultural awareness of large LMs (e.g. GPT-4o) in comparison to native speakers and web content. We observe cultural awareness gaps among LMs. For example, the more Chinese-centric Qwen2.5-72B excels on Chinese cultural entities, literacy, and social norms, while GPT-4o and Mistral-Large match or outperform it on opinion and commonsense about Chinese culture, likely due to differences in their training corpora. To contextualize these gaps, we analyze web content across languages. For questions on social norms and cultural commonsense, where native humans exhibit high confidence, Google search in native languages often return fewer relevant answers than in English. This indicates that such knowledge is often unstated in native contexts while more explicitly articulated in foreign languages Yin et al. (2022).
In summary, our key contributions include:
  • We introduce CARE, a multilingual multicultural human preference dataset focusing on Chinese and Arab cultures (§3).
  • We show that cultural alignment with CARE improves LMs across model families (§4).
  • Motivated by LM performance gaps, we analyze web content across languages and reveal documentation disparities (§5).
Figure 2: Example questions in Arabic and Chinese (with English translations) from the 5 cultural context categories in CARE Refer to caption: Cultural Entities, Cultural Opinion, Social Norms, Cultural Commonsense, and Literacy3.1).
2 Related Work
The widespread use of LMs has sparked research interest in their relevance to diverse cultures Adilazuarda et al. (2024); Shen et al. (2024); Pawar et al. (2024); Liu et al. (2024). Several studies investigate the cultural alignment of LMs by examining their responses to questions from social surveys that reflect human values and attitudes Haerpfer et al. (2021); Cao et al. (2023). It has been consistently shown that LMs favor answers associated with Western culture AlKhamissi et al. (2024); Abdulhai et al. (2023), even when prompted in different languages Masoud et al. (2023); Wang et al. (2023) or after undergoing preference optimization Ryan et al. (2024). While these survey-based studies assess general beliefs on broad cultural topics such as democracy or family values, they have limited coverage of detailed cultural nuances and phenomena such as local customs and daily practices.
Another line of work has developed culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al. (2022); Zhou et al. (2024b), entity-centric cultural benchmarks Naous et al. (2024), and user self-reported cultural experiences Shi et al. (2024). Other works have also constructed QA datasets for different cultural aspects such as culinary customs Palta and Rudinger (2023), norms Rao et al. (2024); Zhan et al. (2024), social etiquette Chiu et al. (2024a, b), and more Arora et al. (2024); Mousi et al. (2024). However, these are mainly designed for evaluating LMs, often in multi-choice QA or text-infilling formats, and are not readily suitable for improving LMs via human preference learning, which requires free-text QA and human ratings on model outputs.
Several multilingual instruction-tuning Muennighoff et al. (2022); Singh et al. (2024); Alyafeai et al. (2024) and preference resources Lian et al. (2023); Bartolome et al. (2023); Ahmadian et al. (2024) have been developed to improve multilingual performance. While they include culturally relevant samples, they primarily offer general instructions and universal preferences for safety and helpfulness. Few works Li et al. (2024a, b); Kirk et al. (2024) attempt to tune LMs on culturally relevant instructions, but they rely on synthesizing data with existing models and mainly express knowledge in English (more on this in §3.1). This may exacerbate biases in the pre-training data and miss authentic viewpoints. Unlike previous studies, our work focuses on culturally aligning LMs by learning from native speakers’ cultural preferences.
3 Constructing CARE [Uncaptioned image]
To study preference optimization of LMs for cultural awareness, we introduce CARE, which consists of 2,580 culture-specific questions and 24.1k responses (one human-written and nine LM-generated per question) with human ratings on a 10-point Likert scale. We begin by identifying and addressing the limitations in existing data resources (§3.1), and then further collect new data for cultural alignment (§3.2).
3.1 Limitations of Existing Data Resources
We start with the Aya dataset Singh et al. (2024), the largest multilingual instruction-tuning resource that contains human-written questions and answers. Though its natural instances collected from fluent native speakers offer authentic viewpoints, only part of its contents focus on culture, and the data quality varies for different languages. After manual inspection and filtering out generic questions (e.g. “How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) question and answer pairs that are culturally relevant in Chinese but only 457 (out of 14,210) samples in Arabic. To expand the Arabic set, we apply the same filtering process to around 2,000 samples from CIDAR Alyafeai et al. (2024), a human-written Arabic instruction dataset, resulting in 162 relevant samples.
Figure 3: Overall coverage of our CARE benchmark per cultural category. CARE provides 16.6$\times$ more coverage of social norm and cultural commonsense questions compared to two existing instruction tuning datasets and four cultural knowledge bases (§3.1).
[…]
Moreover, our analysis reveals that existing resources provide highly unbalanced coverage of various cultural aspects, with a notable lack of data on social norms and commonsense, as shown in Figure 3. We manually classify all questions into one of five categories: (1) Cultural Entities, where the question asks about culture-specific entities Naous and Xu (2025), (2) Cultural Opinion, where the question asks about a subjective interpretation for a cultural entity, (3) Social Norms, where the question is about social interactions between individuals Huang and Yang (2023), (4) Cultural Commonsense, where the question is about daily phenomena that locals may take for granted Shen et al. (2024), and (5) Literacy, where the question is about language usage, proverbs, or slang Wuraola et al. (2024). Figure 2 shows examples for these cultural aspects. Inter-annotator agreement is 0.90, measured by Cohen’s Kappa, for this classification.
Finally, aligning LMs using preference optimization algorithms requires human judgments on both suitable and unsuitable responses, beyond only the existing ground-truth demonstrations.
3.2 Multilingual Cultural Alignment Data
To address the aforementioned shortcomings, we collect new non-English data for cultural alignment as follows:
Social Norm and Commonsense.
To expand the samples on social norms and cultural commonsense, we ask native Chinese and Arabic speakers who are international college students undergoing culture transfer experiences themselves to curate such samples manually. To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad, culture shock, first time to the US, etc.”. Taking inspiration from such discussions, the speakers create 190 samples in Chinese and 196 samples in Arabic. This human-curated data is more authentic than synthetic data, preventing possible inaccuracy and overgeneralization (e.g. “In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
Culture-specific Human Preference.
Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment of LMs through preference optimization. To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs for each question within CARE. We specifically use the instruct version of recent multilingual LMs of Llama3.3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024), Gemma2-27B Team et al. (2024), Mistral-Large, and GPT-4o. We also use their smaller-sized Llama3.1-8B, Qwen2.5-7B, Gemma2-9B, and Mistral-7B.
We then instruct the annotators to rank the generated responses for each question from the most to least culturally appropriate and assign each response a rating on a scale of 1 to 10 (1: poor $\rightarrow$ 10: excellent). Those ratings reflect how well responses match the cultural expectations of native speakers, which we use to construct preference pairs for cultural alignment in our experiments (§4.1). For responses that deviate from answering the questions (e.g., generating similar questions or repetitive content), annotators do not assign a rating.
Human Annotation.
We hire two native Chinese speakers and two native Arabic speakers who are college students to collect and annotate all the data in CARE, as described in §3.1 and §3.2. Besides the 5-class cultural categories, we also annotate each sample with the Associated Culture for experiments in §4.5: Native (questions about the native culture; i.e. Chinese or Arab), Foreign (questions about other cultures), or General (questions that are not specific to a particular culture). For further fine-grained regional cultural knowledge evaluation (Appendix D.1), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide). See Appendix A.1 for the annotation guideline.
Data Split.
For the experiments in §4 and §5, we construct culture-specific test sets with 150 questions each for Chinese and Arab cultures, randomly sampling 30 questions from each of the five cultural categories. The remaining 1,513 Chinese and 644 Arabic questions, along with their responses in CARE, are used to form the training set. This training data also includes questions about other foreign cultures beyond Chinese and Arab.
Chinese Arabic Model Entities Opinion Norms C. sense Literacy Average Entities Opinion Norms C. sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 [Uncaptioned image] [Uncaptioned image]
Table 1: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese and Arab culture in Arabic. Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image]. For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural alignment (colored plot). “C. sense” stands for “Commonsense”.
[…]
Table 2: Performance comparison w.r.t. different preference learning algorithms on CARE[Uncaptioned image] data. Results show the average score improvements over the vanilla model.
4 Aligning LMs for Cultural Awareness
Enabled by CARE, we investigate whether a small amount of high-quality cultural alignment data can improve existing models through preference optimization. We first conduct cultural alignment on medium-sized LMs and show how much improvement we can achieve with CARE in §4.1 and §4.2. Next, we examine the impact of scaling data and model size (§4.3) and whether cultural alignment sacrifices LMs’ capabilities on general tasks (§4.4). Finally, we analyze the importance of including data from diverse cultures during alignment (§4.5).
4.1 Experiment Setup
We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE on up to 8 Nvidia A40 GPUs.
Human Preference Optimization.
We experiment with different preference learning algorithms, including DPO Rafailov et al. (2024), KTO Ethayarajh et al. (2024), and SimPO Meng et al. (2024). These algorithms require preference pairs that contrast two responses for each question: one culturally appropriate and one that is less appropriate. We use the highest- and lowest-scored responses from the human preference data collected in CARE. We find that all three offline preference algorithms perform similarly well, as shown in Table 3.2; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.
Baselines.
First, we analyze the effectiveness of learning from cultural preference compared to general preference. We align LMs using existing human preference datasets, such as OpenOrca Lian et al. (2023) and UltraChat Bartolome et al. (2023), which focus on truthfulness, honesty, and helpfulness. For a fair comparison, we randomly sample an equal number of training pairs from these datasets as in CARE’s training set. Second, we examine the effectiveness of SFT on CARE versus Alpaca Cui et al. (2023); Chen et al. (2023), a general instruction-tuning dataset with Chinese and Arabic adaptations. Third, for prompting-based approaches, we explore CoT prompting Wei et al. (2022) and role-play prompting Kong et al. (2023). See prompts used in Appendix B.
LM-as-a-judge Evaluation and its Reliability.
For evaluation, we adopt the LM-as-a-judge Zheng et al. (2023) strategy where we prompt a powerful LM to evaluate how aligned the generated response is to the human reference response. We use GPT-4o as the LM judge and prompt it to provide reasoning and a score on a scale of 1 to 10. We validate this setup by comparing it to the ratings of the native-speaking human annotators, where we achieve a high Pearson correlation of 0.93 between human and GPT-4o ratings. More details and our evaluation prompts can be found in Appendix C.
[…]
Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese and Arab cultures for a variety of prompting approaches, fine-tuning, and alignment using cultural (CARE) vs. generic instruction-tuning (Multilingual Alpaca) and preference (OpenOrca/UltraChat) data. SFT is performed on the instruction data only, while alignment conducts DPO on the human preference pairs. Aligning with cultural preferences enhances cultural awareness more effectively than using general preferences, with this improvement remaining consistent across various prompting setups.
4.2 Main Results
Table 1 presents the results of medium-sized LMs before and after cultural alignment using CARE, while Table 3 compares these results with various baselines. We have the following key findings:
Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) in both Chinese and Arabic compared to the vanilla checkpoints, even for these already instruction-tuned models. This resonates with the findings of Zhou et al. (2024a) and shows that a relatively small amount of carefully curated data by humans can improve LMs’ alignment. We also see a noticeable gap among LMs developed by different regions. The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) overall by aligning with CARE. Interestingly, the aligned version of Gemma2-9B outperforms Qwen2.5-7B on social norms and commonsense for Chinese (more on this in §5.2). We also test its generalization capabilities on other culture-related tasks in Appendix D.
Cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across all five cultural categories after alignment. However, alignment shows limited gains where the model’s initial performance is poor. Mistral-7B is not improved on Chinese entities and literacy, where its starting scores are only 3.03 and 2.43. Similarly, Qwen2.5-7B, does not benefit from alignment on Arabic entities. This suggests that base models need a foundational level of cultural knowledge for alignment to be effective.
Culture-specific human preference data is helpful beyond general human preference. Table 3 shows that alignment with CARE’s cultural preference consistently outperforms general preference datasets by up to 42%. Additionally, cultural alignment with CARE brings improvement in both CoT and role-play prompting. While culture-specific SFT outperforms generic instruction tuning, alignment yields greater gains, highlighting the importance of human cultural preference data over ground-truth demonstrations alone.
4.3 How does scaling affect alignment?
We examine how scaling both model size and data volume influences the performance before and after cultural alignment. Specifically, we consider the Qwen2.5 {1.5B, 3B, 7B, 14B} and the Gemma2 {2B, 9B} series, which offer multiple model sizes. For data sizes, we align models using DPO on different proportions {0%, 25%, 50%, 75%, 100%} of Chinese cultural preference pairs within CARE.
As shown in Figure 4, consistent improvements are observed across different model and data sizes. Even a relatively small amount (e.g., 25%) of data in CARE can lead to improvements, particularly for smaller-sized LMs. Scaling up the data leads to progressively better performance, highlighting the potential benefits of employing more cultural preference data in future work.
Figure 4: Impact of model size and preference data volume on cultural awareness performance. The average scores (1: poor $\rightarrow$ 10: excellent) of aligned models are plotted against different % of preference pairs in CARE. Improvements are achieved across different model sizes and data sizes, in comparison to the vanilla model (0%).
4.4 Does alignment impact general capability?
We then examine whether cultural alignment impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, and English: ArabicMMLU Koto et al. (2024), ChineseMMLU Li et al. (2023), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), and WinoGender Rudinger et al. (2018). We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraChat and OpenOrca datasets (totaling 3,718 samples), the culturally aligned LM with cultural human preference from CARE, and the aligned LM with both general and CARE’s cultural preference.
The results in Figure 5 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural alignment does not hinder the model’s overall capabilities.
Figure 5: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese and Arabic) and English. Slight differences are observed among the LM variants after applying different alignment setups.
4.5 Should we align with only native data?
We analyze the impact of incorporating data about different cultures in the alignment process. To ensure a fair comparison, we select an equal number of samples for each trial. Specifically, we align Llama3.1-8B-Instruct with 676 Chinese and 546 Arabic pairs sampled from three different contexts in CARE: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign).
Results are shown in Figure 6. We find that performance drops on native culture when the model is aligned using only foreign culture samples, but improves when aligned with the same native culture samples (4.69 $\rightarrow$ 4.87 in Chinese and 2.98 $\rightarrow$ 3.50 in Arabic); interestingly, performance is further improved when the model is aligned by combining native and foreign cultural data. This indicates that expanding geographical coverage in alignment could further enhance cultural awareness.
[…]
Table 4: Average scores on CARE of larger LMs and web content, evaluated by the judge LM across all samples (1: poor $\rightarrow$ 10: excellent). “Human Awareness (%)” indicates the percentage of questions for which natives know the correct answer. Qwen2.5 developed by Chinese-based Alibaba excels on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense. For social norms and cultural commonsense, where native humans show high confidence, Google search often returns fewer relevant answers.
5 Prompting LMs for Cultural Awareness
Figure 6: Average scores of Llama3.1-8B-Instruct on native and foreign cultures when aligned using data from native, foreign, or mixed cultures. The highest performance on native culture is achieved when mixing native and foreign samples during alignment.
Beyond improving medium-sized LMs with culture-specific human preference, we evaluate the performance of larger LMs and compare with native human familiarity and web-based sources (§5.1). To better understand these differences, we examine the role of querying language (§5.2).
5.1 Main Results
We follow the evaluation setup in §4.1 and assess the larger ($>$ 25B parameters) versions of the evaluated LMs with zero-shot prompting. Table 4 shows the results. Larger models are better for both Chinese and Arabic, with GPT-4o achieving the best overall performance. The more Chinese-centric model Qwen2.5-72B excels in the knowledge about cultural entities (8.42), social norms (9.06), and literacy (8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture. This is interesting and likely because cultural commonsense knowledge is often unstated in the native language, while more balanced and thoughtful opinions may be expressed in foreign languages. To illustrate and test this hypothesis, we conduct further analysis in §5.2.
Human Awareness.
To assess human familiarity with the culture-specific questions, we ask native-speaking annotators whether they knew the correct answers without looking them up when presented with questions from CARE. Table 4 shows the percentages of questions humans could answer immediately with confidence, which are very high for social norms and commonsense knowledge.
Search Engine.
To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google [^1] Google programmable search engine API: https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE. Results are shown in Table 4. Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.
(a) Average scores achieved by Qwen2.5-72B-Instruct and GPT-4o when prompted in native languages versus English.
(b) Average scores of retrieved webpage main content when searching CARE questions in native languages versus English.
Figure 7: Performance comparison of cultural awareness in different languages. Comparison between (a) LM performance and (b) retrieved webpage content, showing that native-language questions capture cultural nuances for literacy, opinion, and entities. However, the gaps narrow or sometimes reverse for cultural norms and commonsense.
5.2 Cross-lingual Analysis
We translate all questions and responses from CARE that are written in Chinese and Arabic into English to enable cross-lingual evaluation.
Figure 7(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than for Arabic. This advantage is also especially obvious for the literacy category, while the gap narrows in categories such as cultural commonsense. A possible explanation is that English corpora contain fewer discussions about Chinese or Arab cultures.
We also examine how language choice impacts search engine results. As shown in Figure 7(b), cultural information on entities, opinions, and literacy tends to be of higher quality when retrieved in native languages. Interestingly, this trend reverses for questions related to cultural commonsense and social norms, such as “In China, does the ticket time indicate when the feature movie starts, or are there trailers played before the main feature?”, where English searches yield higher-quality content. This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, whereas non-native speakers are more likely to actively seek out and document it. A similar pattern is observed for Arabic. For a small number of Arabic questions (10% opinion, 14% social norms, 14% commonsense, and 7% literacy), the search engine also failed to return any results. One such example is “A Jordanian friend mentioned that it’s polite not to open a gift in front of the giver. Is this true?”.
6 Conclusion
We introduce CARE, a human-written, multilingual resource comprising 2,580 culture-specific questions about Chinese and Arab cultures, along with responses that include human preferences. Through extensive experiments, we investigate that a small amount of high-quality cultural alignment data can improve existing LMs via preference optimization, across different model families and scales. This cultural-specific alignment offers benefits that go beyond generic human preference and proves effective under various prompting setups. Our analyses also reveal gaps in cultural awareness among LMs, native-speaking humans, and search engines. For example, cultural commonsense is often left implicit in the native language but is more explicitly mentioned in foreign-language content. By releasing CARE, we hope to encourage the development of more inclusive and culturally adaptive LMs.
Limitations
In this work, as we investigate how to improve LMs by aligning with multilingual multicultural human preference data, we also identified a couple of areas for further exploration. First, while filtering the Arabic portion of the Aya dataset, we notice a large number (3,831) of questions centered on Islamic knowledge and religious history. We exclude them, as religion-specific understanding in LMs falls outside the scope of our study that focuses on culture-specific phenomena and factual knowledge, as defined in §3.1. Nonetheless, future research on religious QA or alignment could adopt a similar approach to CARE by collecting preference data from credible religious sources.
Second, LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding. For instance, when asked “In China, what does the term “laowai” mean when referring to foreigners?”, LMs often incorrectly interpret the term as disrespectful. In reality, “laowai” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations. While CARE includes examples that clarify such misunderstandings through human-annotated responses, future studies in cultural red-teaming can further investigate and address these failure modes.
Acknowledgments
The authors would like to thank Nour Allah El Senary, Jad Matthew Bardawil, Chao Jiang, and Yuming Pan for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback. We would also like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This work was funded in part through a Sony Faculty Innovation Award, and by the NSF under grant number IIS-2144493, IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A CARE: Details
A.1 Annotation Details
To construct the question-answer pairs in CARE, we recruited two native Chinese speakers and two native Arabic speakers who are undergraduate students. The annotators were paid at the rate of 16 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs. The detailed annotation guideline for dataset construction is provided in Figures 16 to 20, and the instruction for preference annotation is provided in Figures 14. We also show our rank-and-rate web interface for collecting human cultural preference pairs in Figure 15.
A.2 Statistics
Table 5 shows the detailed statistics for Chinese and Arab culture in CARE across the cultural categories, stratified by the data source. The questions specific to Chinese culture are written in Chinese, while the questions specific to Arab culture are written in Arabic. The culturally-relevant samples obtained from the instruction-tuning dataset mostly fall within the cultural entities, cultural opinion, or literacy categories, while samples obtained from cultural knowledge bases provide more social norms and cultural commonsense data.
Table 6 shows the detailed statistics for other foreign cultures in CARE for each cultural category. These samples were obtained when filtering the instruction tuning datasets and cultural knowledge bases (§3.1). We wrote these foreign samples in both Chinese and Arabic versions and used them as part of our training set and in our analysis on the impact of the source culture in alignment (§4.5). We also manually collected more samples for social norms and cultural commonsense from online websites to cross 100 samples in each category. We ensured these foreign samples do not include anything relevant to Chinese or Arab cultures.
Table 7 shows the detailed statistics for Chinese and Arabic samples in CARE across the geographic scope. Most samples fall into nationwide category, while sub-nationwide samples provide insights about more detailed cultural information, continent-wide and worldwide samples assess understanding between several cultures.
Appendix B Implementation Details
DPO Alignment.
We perform preference optimization using DPO for 3 epochs until the loss converges. Training is done with a batch size of 128 on 4-8 NVIDIA A40 GPUs. We tune the learning rate in the range $\{1e^{-7},5e^{-7},9e^{-7}\}$ and set beta for DPO as 0.1. The training involves full fine-tuning with 5 warmup steps and employs a linear learning rate scheduler.
LM Inference.
For open-sourced LMs, we run inference on 1 NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al. (2023). We perform decoding by setting the following parameters {temperature=0.7, top_p=1}. We limit the context length by setting {max_model_len=2048}. We also limit the number of generated tokens by the models by setting {max_tokens=1024}. For the closed-source GPT-4o LM, we run inference with Azure OpenAI API.
Baselines Prompt Templates.
The prompt templates used for our prompting-based baselines (§4.2) are provided in Figure 8 (role-play prompting) and Figure 9 (CoT prompting).
Role-play Prompt Template
You are a native [Chinese/Arab] person, familiar with
[Chinese/Arab] culture and traditions.
{Question}
Figure 8: Prompt template for role-play inference. The LM is told to take on the persona of a native [Chinese/Arab] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
CoT Prompt Template
You are a helpful assistant.
{Question}
Let’s think step by step.
Figure 9: Prompt template for CoT inference. We provide the LM with the test question in {Question}, then ask it to think step by step when providing the answer.
Chinese Arabic Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92
Table 8: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both native culture and foreign culture.
Appendix C LM-as-a-Judge
C.1 Evaluation Prompts
We instruct GPT-4o as the judge LM to score a model’s response to culture-specific questions in CARE. For each cultural category, we provide the judge LM with a detailed evaluation guideline, the culture-specific question, the generated response, and the human reference response, and ask it to score the response on the 1-10 scale. Our evaluation prompt templates are provided in Figure 11 (Entities & Opinion), Figure 12 (Norms & Commonsense), and Figure 13 (Literacy).
C.2 Correlation with Human Ratings
Fig 10 shows the rating distribution of GPT-4o and native speakers. We can see a clear correlation between both rating distributions. We also calculate correlation metrics and obtain a Pearson correlation of 0.933, Spearman correlation of 0.901, and Kendall’s Tau correlation of 0.733, all indicating high agreement between the judge LM ratings and human ratings. One difference we notice is that humans tend not to assign extreme ratings (1 or 10), yet the LM judge has a higher frequency of those ratings. This is mostly noticeable at the 9-10 rating range where the LM ratings are more equally distributed between a rating of 9 and 10, but the human ratings mostly consist of 9.
Figure 10: Rating score distributions of the judge LM (GPT-4o) and native human evaluators. The judge LM highly correlates with human ratings.
Appendix D Additional Results
D.1 Performance by Geographic Scope
Table 8 shows the results achieved by different LMs when stratified by geographic scope. We find that the models struggle the most with sub-nationwide questions. Most models achieve the best scores on continent-wide or worldwide questions.
D.2 Performance on other cultural tasks
We also evaluate the performance of LMs after cultural alignment with CARE on other cultural tasks using out-of-domain culture-related data. Specifically, we use the Include benchmark Romanou et al. (2025), as it offers multilingual versions for each culture and employs a multi-choice format, enabling us to assess the robustness of our alignment. From Table 9, we observe that alignment with culture-specific human preference can also benefit the performance on the Include benchmark, indicating the generalization capability of CARE.
[Uncaptioned image] Gemma2-9B [Uncaptioned image] Llama3.1-8B [Uncaptioned image] Qwen2.5-7B [Uncaptioned image] Mistral-7B Vanilla CARE Vanilla CARE Vanilla CARE Vanilla CARE Include (zh) 57.79 59.44 51.00 54.68 75.22 75.59 40.36 40.73 Include (ar) 57.42 54.71 50.54 51.08 56.15 55.97 35.32 36.95
Table 9: Accuracy (%) on the Include benchmark for Chinese and Arabic subsets, before and after cultural alignment with CARE. Improvements in the Include benchmark can also be observed after alignment with CARE.
Appendix E Responsible NLP Checklist
In our work, we ensure that the use of existing artifacts is consistent with their intended use. Specifically, all samples collected for the CARE dataset are sourced from open-sourced datasets, adhering strictly to the original access conditions intended for research purposes. We meticulously filter the data to remove any instances of toxic or offensive language and ensure that it does not contain any personally identifiable information. For the artifacts we create, we clearly specify their intended use within research contexts, maintaining compatibility with the original access conditions. We acknowledge the use of GPT-4 for grammar checking and word polishing.
Cultural entities and opinion evaluation
Please serve as an unbiased evaluator and critically assess the quality of the assistant’s response to the user’s question
presented below. When evaluating, focus on the following aspects:
1. **Accuracy**: Is information in the response factually correct and up-to-date?
2. **Depth**: Does the response demonstrate a deep understanding of the topic, providing thorough explanations and context?
3. **Relevance**: Does the response stay focused on the specific question without including unnecessary information?
Begin your evaluation with a brief commentary explaining your judgment on each of these criteria. Aim to be as objective as
possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response is highly inaccurate, lacks detail, contains significant incorrect information, and/or
includes irrelevant information.
- **3-4**: Below Average — The response is partially accurate, addresses some parts of the question but lacks detail, and
may include irrelevant information.
- **5-6**: Average — The response is moderately accurate but may contain minor errors, addresses most parts of question with
adequate detail, and is mostly relevant.
- **7-8**: Good — The response is mostly accurate, addresses all parts of the question with good detail, and is relevant
with minimal irrelevant information.
- **9-10**: Excellent — The response is highly accurate, provides comprehensive detail, and contains no irrelevant information.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Golden answer: {answer}
## Assistant’s response: {response}
Figure 11: LM-as-a-judge prompt template for cultural entities and opinion questions.
Social norms and commonsense evaluation
Please serve as a critical evaluator and rigorously assess the quality of the assistant’s response to the user’s question
shown below. When evaluating, prioritize the following stringent criteria:
1. **Accuracy**: Based on the golden answer, is the information factually correct and representative of real-world cultural
context?
2. **Comprehensiveness**: Does the response provide clear explanations and consider possible nuances or exceptions within
the cultural context?
3. **Relevance**: Is the response perfectly tailored to the specific cultural context, without any generalizations or
inaccuracies?
Begin your evaluation with a detailed commentary critically analyzing each of these criteria. Strive to be as objective
and discerning as possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response fails to meet basic expectations for accuracy or relevance, showing major misunderstandings or
errors.
- **3-4**: Below Average — The response has substantial inaccuracies or omissions, only partially addressing the user’s needs.
- **5-6**: Average — The response is fairly accurate and relevant but lacks depth, missing important details or subtleties.
- **7-8**: Good — The response is accurate and covers most aspects well, though it may lack in minor details or perfect
contextual alignment.
- **9-10**: Excellent — The response is outstanding in all respects; it is precise, detailed, fully relevant, and excellently
contextualized.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Golden Answer: {answer}
## Assistant’s response: {response}
Figure 12: LM-as-a-judge prompt template for social norms and commonsense questions.
Literacy evaluation
Please serve as a critical evaluator and rigorously assess the quality of the assistant’s response to the user’s question shown
below. When evaluating, prioritize the following stringent criteria:
1. **Accuracy**: Is the information in the response factually correct and contextually appropriate?
2. **Interpretation**: Does the response offer insightful and well-supported interpretations of the literary work or topic?
3. **Textual Evidence**: Does the response appropriately reference and analyze specific parts of the text to support its points
when necessary?
4. **Relevance**: Does the response stay focused on specific question without including unnecessary information?
Begin your evaluation with a detailed commentary critically analyzing each of these criteria. Strive to be as objective and
discerning as possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response fails to meet basic expectations for accuracy or relevance, showing major misunderstandings or
errors.
- **3-4**: Below Average — The response has substantial inaccuracies or omissions, only partially addressing the user’s needs.
- **5-6**: Average — The response is fairly accurate and relevant but lacks depth, missing important details or subtleties.
- **7-8**: Good — The response is accurate and covers most aspects well, though it may lack in minor details or perfect contextual
alignment.
- **9-10**: Excellent — The response is outstanding in all respects; it is precise, detailed, fully relevant, and excellently
contextualized.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Reference Answer: {answer}
## Assistant’s response: {response}
Figure 13: LM-as-a-judge prompt template for literacy questions.
Figure 14: Cultural preference annotation instructions.
Figure 15: The interface for annotating culture-specific human preference.
Figure 16: Annotation Guideline (1/5).
Figure 17: Annotation Guideline (2/5).
Figure 18: Annotation Guideline (3/5).
Figure 19: Annotation Guideline (4/5).
Figure 20: Annotation Guideline (5/5).

### v4

@@
 Abstract
-Existing language models (LMs) often exhibit a Western-centric bias and struggle to represent diverse cultural knowledge.
-Previous attempts to address this rely on synthetic data and express cultural knowledge only in English.
-In this work, we study whether a small amount of human-written, multilingual cultural preference data can improve LMs across various model families and sizes.
-We first introduce CARE, a multilingual resource of 24.1k responses with human preferences on 2,580 questions about Chinese and Arab cultures, all carefully annotated by native speakers and offering more balanced coverage.
-Using CARE, we demonstrate that cultural alignment improves existing LMs beyond generic resources without compromising general capabilities.
-Moreover, we evaluate the cultural awareness of LMs, native speakers, and retrieved web content when queried in different languages.
-Our experiment reveals regional disparities among LMs, which may also be reflected in the documentation gap: native speakers often take everyday cultural commonsense and social norms for granted, while non-natives are more likely to seek out and document them actively.
-CARE is publicly available at https://github.com/Guochry/CARE (we plan to add Japanese data in the near future).
-CARE: Aligning Language Models for Regional Cultural Awareness
+Language Models (LMs) are typically tuned with human preferences to produce helpful responses, but the impact of preference tuning on the ability to handle culturally diverse queries remains understudied.
+In this paper, we systematically analyze how native human cultural preferences can be incorporated into the preference learning process to train more culturally aware LMs.
+We introduce CARE, a multilingual resource containing 3,490 culturally specific questions and 31.7k responses with native judgments.
+We demonstrate how a modest amount of high-quality native preferences improves cultural awareness across various LMs, outperforming larger generic preference data.
+Our analyses reveal that models with stronger initial cultural performance benefit more from alignment, leading to gaps among models developed in different regions with varying access to culturally relevant data.
+CARE will be made publicly available at https://github.com/Guochry/CARE.
+CARE: Assessing the Impact of Multilingual Human Preference Learning on Cultural Awareness
 […]
+Figure 1: Example LM responses to culture-specific questions in the native languages.
+The base LM (Refer to caption) Llama3.1-8B fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption), Arab (Refer to caption), and Japanese (Refer to caption) cultures.
 1 Introduction
-The impressive capabilities of multilingual language models (LMs) lead to their widespread use across various cultural groups Ouyang et al.
-(2022); Dubey et al.
+After large-scale pre-training, Language Models (LMs) typically undergo a crucial post-training phase (aka “alignment”), which includes supervised fine-tuning and preference tuning, to improve their ability to follow instructions and align with human preferences Ouyang et al.
+(2022); Rafailov et al.
 (2024).
-However, many studies have shown that LMs often lack knowledge of diverse cultures and are biased towards Western culture Naous et al.
-(2023); AlKhamissi et al.
-(2024); Naous and Xu (2025).
-While much focus has been on demonstrating these cultural biases, less effort Li et al.
-(2024a, b); Kirk et al.
-(2024) has been made to address them.
-Existing works have largely been limited to expressing cultural knowledge in only English or generating synthetic data with existing LMs for fine-tuning.
-Figure 1: Example LM responses to culture-specific questions in Chinese and Arabic.
-The base LM (Refer to caption) Llama3.1-8B-Instruct fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption) and Arab (Refer to caption) cultures.
-In this work, we explore the benefits of learning from multilingual multicultural human preference data, and demonstrate how even a modest dataset of 2.6k high-quality human-written samples can significantly enhance LM’s cultural awareness across various model families and sizes.
-Specifically, we first introduce the Cultural Alignment REsource (CARE [Uncaptioned image]), comprising 24.1k responses (one human-written and nine LM-generated per question) with native human preference using a 10-point scale on 2580 multilingual questions focusing on Chinese and Arab cultures (825 and 697 in respective languages; additional 1,058 covering other cultures).
-Compared with existing datasets Yin et al.
-(2022); Palta and Rudinger (2023); Shi et al.
-(2024); Chiu et al.
-(2024b), our data offers more balanced coverage and authentic viewpoints, especially for topics such as social norms and cultural commonsense, where previous works have limited data coverage.
-To achieve this, we have native speakers both re-annotate existing data resources, and curate new data on those rarely documented yet deeply embedded aspects of daily life.
-Using CARE, we demonstrate that multilingual human cultural preference enhances LMs’ cultural awareness through alignment strategies, such as direct preference optimization (DPO, Rafailov et al., 2024), without compromising general capabilities.
-This cultural alignment proves beneficial beyond standard human preference datasets, with consistent gains across both chain-of-thought (CoT) and role-play prompting settings.
-We further analyze post-alignment performance across different LMs and find that for cultural alignment to be effective, base LMs need a foundational level of cultural knowledge.
-That is, cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs.
-We then examine the impact of cross-cultural alignment, and notice that aligning models with data from multiple cultures results in greater improvements than using single-culture data.
-This highlights the benefits of expanding geographical coverage in LM alignment.
-Moreover, we evaluate the cultural awareness of large LMs (e.g.
-GPT-4o) in comparison to native speakers and web content.
-We observe cultural awareness gaps among LMs.
-For example, the more Chinese-centric Qwen2.5-72B excels on Chinese cultural entities, literacy, and social norms, while GPT-4o and Mistral-Large match or outperform it on opinion and commonsense about Chinese culture, likely due to differences in their training corpora.
-To contextualize these gaps, we analyze web content across languages.
-For questions on social norms and cultural commonsense, where native humans exhibit high confidence, Google search in native languages often return fewer relevant answers than in English.
-This indicates that such knowledge is often unstated in native contexts while more explicitly articulated in foreign languages Yin et al.
+However, the effects of alignment on the cultural awareness of multilingual LMs (i.e., how well they understand and generate appropriate responses to diverse, culturally relevant queries) remains understudied, partly because frontier open-weight models with multilingual support (e.g., Llama-3, Qwen-2.5) have become available very recently around mid-2024.
+Researchers have only examined the off-the-shelf aligned LMs (e.g., Aya, Tulu 2) Naous et al.
+(2024); Ryan et al.
+(2024), while others explored instruction fine-tuning with culturally relevant data Li et al.
+(2024a).
+Dataset Culture- Topics Manually- Human Languages Format
+Specific Curated Preferences
+Cultural Evaluation
+     Fork (Palta and Rudinger) food-related customs en multiple-choice
+     CulturalBench (Chiu et al.) daily life, social etiquette, wider society en multiple-choice
+     GeoMLAMA (Yin et al.) cultural commonsense en, fa, hi, sw, zh multiple-choice
+     BLEnD (Myung et al.) food, sports, family, education, holidays, work-life 13 languages multiple-choice, free text
+     Include (Romanou et al.) general knowledge, social science, professional certifications, etc. 44 languages multiple-choice
+     CAMeL (Naous et al.) beverage, clothing, food, location, religion, sports, etc. ar masked prompts
+     CANDLE (Nguyen et al.) geography, religion, occupation, food, clothing, etc. en assertions
+     LLM-GLOBE (Karinshak et al.) cultural values zh, en multiple-choice, free text
+     CulturalTeaming (Chiu et al.) red-teaming cultural questions en multiple-choice
+     SHADES (Mitchell et al.) culture-specific stereotypes 16 languages stereotyped statements
+     JMMMU (Onohara et al.) art, heritage, history, business, science, medicine, etc. ja multiple-choice
+[…]
+     Aya (Singh et al.) [Uncaptioned image] news, stories, recipes, scientific texts, etc. 65 languages free text
+     CIDAR (Alyafeai et al.) [Uncaptioned image] technology, translation, poetry, grammar, etc. ar free text
+     Cameleval (Qian et al.) information provision, reasoning, creative writing, etc. ar free text
+     PRISM (Kirk et al.) cross-cultural controversies en multi-turn conversations
+     Palm (Alwajih et al.) history, celebrations, sports, literature, etc. ar free text
+     CULTUREINSTRUCT (Pham et al.) art, cuisine, cultural norms, festivals, history, etc. en free text
+     CultureBank (Shi et al.) social norms, food, communication, festivals, etc. en free text
+     CulturePark (Li et al.) human belief, norm, custom en free text
+     CultureLLM (Li et al.) value survey 9 languages multiple-choice
+CARE (Ours) [Uncaptioned image] cultural entities, opinion, norms, commonsense, literacy ar, ja, zh free text
+Table 1: Comparison of datasets for studying LMs’ cultural awareness.
+CARE is a multilingual, human-annotated preference dataset specifically grounded in culture.
+([Uncaptioned image]) indicates resources that include some cultural considerations (in the form of culture-related questions or by recruiting native annotators), while cultural coverage is not their only or primary focus (see § 3.1).
+Representative examples from each dataset are provided in Appendix A.4.
+In this paper, we systematically analyze the extent to which and under what conditions preference optimization (e.g., DPO, KTO, SimPO) can help LMs align with regional user expectations along five key dimensions: cultural entities, cultural commonsense, social norms, opinions, and literacy (see examples in Figure 1).
+To enable our study, we create [Uncaptioned image] CARE, a multilingual culture-specific dataset with human preference judgments on 31.7k LLM/Human-written responses to 3,490 questions about Chinese, Arab, and Japanese cultures.
+Questions are drawn from a variety of instruction tuning datasets Singh et al.
+(2024); Alyafeai et al.
+(2024) and evaluation resources Chiu et al.
+(2024b); Palta and Rudinger (2023) that we manually identified as culturally-relevant.
+We further collected additional questions to improve the coverage on regionally popular but less digitally documented commonsense knowledge and social norms, with help from native speakers who have overseas experience.
+Leveraging CARE, we show that a modest, high-quality set of native preferences can enhance LMs’ cultural awareness across model families and sizes, outperforming generic preference data despite having 3–17$\times$ fewer samples, and generalizing to out-of-domain cultural tasks Myung et al.
+(2024); Romanou et al.
 (2025).
-In summary, our key contributions include:
-  • We introduce CARE, a multilingual multicultural human preference dataset focusing on Chinese and Arab cultures (§3).
-  • We show that cultural alignment with CARE improves LMs across model families (§4).
-  • Motivated by LM performance gaps, we analyze web content across languages and reveal documentation disparities (§5).
-Figure 2: Example questions in Arabic and Chinese (with English translations) from the 5 cultural context categories in CARE Refer to caption: Cultural Entities, Cultural Opinion, Social Norms, Cultural Commonsense, and Literacy3.1).
+We then conduct controlled studies on various LMs and uncover several important insights for cultural alignment: (1) Base model strength matters: aligning base models with stronger initial cultural awareness such as Qwen2.5-7B and Gemma2-9B lifts scores across all cultural categories, whereas models with weaker initial performance hardly improve after alignment, consistent with the post-training analyses of Ivison et al.
+(2023).
+(2) Data efficiency scales: even 25% of CARE can lead to a 74% improvement, and performance rises with scaling to the full dataset.
+(3) Cross-culture synergy: mixing samples about local and foreign cultures in preference learning improves overall awareness of local cultures (e.g., Chinese 4.69 $\rightarrow$ 4.87, Arabic 2.98 $\rightarrow$ 3.50).
+Beyond these insights, we observe that performance on different cultural dimensions varies across model families.
+For example, the Chinese-centric Qwen2.5-72B excels on Chinese entities and social norms but lags behind GPT-4o on cultural commonsense.
+This suggests that disparities in data access may lead to uneven cultural competencies across models.
+The querying language also matters: models consistently perform better when queried about culture-specific phenomena in the native language (i.e., Arabic, Chinese, Japanese) rather than English, with the gap narrowing for cultural commonsense.
+A likely explanation is that everyday cultural commonsense is often unspoken in native contexts while more explicitly expressed and asked about in foreign languages Yin et al.
+(2022).
+Our analysis with search engines further supports this, highlighting the importance of broad cultural and linguistic coverage in training data.
 2 Related Work
+Cultural Evaluation.
 The widespread use of LMs has sparked research interest in their relevance to diverse cultures Adilazuarda et al.
@@
 (2024).
-Several studies investigate the cultural alignment of LMs by examining their responses to questions from social surveys that reflect human values and attitudes Haerpfer et al.
+Several studies investigate LMs’ alignment to different cultures by examining their responses to social surveys that reflect human values and attitudes Haerpfer et al.
 (2021); Cao et al.
@@
 (2023); Wang et al.
-(2023) or after undergoing preference optimization Ryan et al.
+(2023); Rystrøm et al.
+(2025) or after preference optimization Ryan et al.
 (2024).
-While these survey-based studies assess general beliefs on broad cultural topics such as democracy or family values, they have limited coverage of detailed cultural nuances and phenomena such as local customs and daily practices.
-Another line of work has developed culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al.
+Another line of work develops culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al.
 (2022); Zhou et al.
@@
 (2024).
-Other works have also constructed QA datasets for different cultural aspects such as culinary customs Palta and Rudinger (2023), norms Rao et al.
+Other works have constructed QA datasets for different cultural aspects, such as culinary customs Palta and Rudinger (2023), norms Rao et al.
 (2024); Zhan et al.
 (2024), social etiquette Chiu et al.
-(2024a, b), and more Arora et al.
+(2024a, b); Qiu et al.
+(2025), and more Arora et al.
 (2024); Mousi et al.
 (2024).
-However, these are mainly designed for evaluating LMs, often in multi-choice QA or text-infilling formats, and are not readily suitable for improving LMs via human preference learning, which requires free-text QA and human ratings on model outputs.
-Several multilingual instruction-tuning Muennighoff et al.
+However, they are primarily designed for evaluation, often in a multiple-choice QA format, and are not well-suited for aligning LMs through preference optimization, which ideally relies on free-text QA and human preference data.
+(see Table 1 and §3.1).
+Cultural Fine-tuning.
+Several resources Muennighoff et al.
 (2022); Singh et al.
 (2024); Alyafeai et al.
-(2024) and preference resources Lian et al.
+(2024); Lian et al.
 (2023); Bartolome et al.
 (2023); Ahmadian et al.
-(2024) have been developed to improve multilingual performance.
-While they include culturally relevant samples, they primarily offer general instructions and universal preferences for safety and helpfulness.
-Few works Li et al.
-(2024a, b); Kirk et al.
-(2024) attempt to tune LMs on culturally relevant instructions, but they rely on synthesizing data with existing models and mainly express knowledge in English (more on this in §3.1).
-This may exacerbate biases in the pre-training data and miss authentic viewpoints.
-Unlike previous studies, our work focuses on culturally aligning LMs by learning from native speakers’ cultural preferences.
+(2024) have been developed to improve LMs’ multilingual performance.
+While they include culturally relevant samples, they primarily offer general instructions and preferences for safety and helpfulness.
+A few other works have investigated cultural adaptation of LMs via fine-tuning strategies Li et al.
+(2024a, b); Shi et al.
+(2024); Kirk et al.
+(2024); Choenni et al.
+(2024); Yuan et al.
+(2024); Pham et al.
+(2025); Alwajih et al.
+(2025); Xu et al.
+(2025); Yao et al.
+(2025).
+Unlike past studies (see Table 1), CARE provides human preferences on culture-specific topics in three languages (not translated from English), enabling a direct study of how multilingual preference optimization improves models’ cultural awareness and by how much.
+Multilingual Preference Optimization
+Existing work aims to improve the alignment of LMs with human preferences across different languages, either by synthesising preference data Pan (2024); Hwang (2024); 2A2I (2024) or developing translation-based methods She et al.
+(2024); Yang et al.
+(2024b).
+Only a few studies have released native multilingual human preference datasets that are neither translated from English nor rated by AI models.
+Notable examples include OpenAssistant Köpf et al.
+(2023) and HelpSteer3-Preference Wang et al.
+(2025b), both of which focus on general-purpose interactions.
+We focus on culturally-relevant preferences for an in-depth analysis of multilingual multicultural alignment.
+Figure 2: Overall coverage per cultural category.
+CARE provides 16.6$\times$ more social norm and commonsense questions compared to 2 instruction tuning datasets (Aya and CIDAR) and 4 cultural knowledge bases (CultureBank, CultureBench, GeoMLAMA, FORK) combined.
 3 Constructing CARE [Uncaptioned image]
-To study preference optimization of LMs for cultural awareness, we introduce CARE, which consists of 2,580 culture-specific questions and 24.1k responses (one human-written and nine LM-generated per question) with human ratings on a 10-point Likert scale.
-We begin by identifying and addressing the limitations in existing data resources (§3.1), and then further collect new data for cultural alignment (§3.2).
+We introduce CARE, a multilingual human preference dataset that consists of 3,490 culture-specific questions and 31.7k human/LLM responses with human preference ratings.
 3.1 Limitations of Existing Data Resources
+In this section, we notice limitations of existing resources and take action to fix them.
 We start with the Aya dataset Singh et al.
 (2024), the largest multilingual instruction-tuning resource that contains human-written questions and answers.
-Though its natural instances collected from fluent native speakers offer authentic viewpoints, only part of its contents focus on culture, and the data quality varies for different languages.
+Though its samples are collected from native speakers, only part of its content focuses on culture, as many examples consist of general questions.
 After manual inspection and filtering out generic questions (e.g.
-“How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) question and answer pairs that are culturally relevant in Chinese but only 457 (out of 14,210) samples in Arabic.
+“How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) and 600 (out of 3,264) question and answer pairs that are culturally relevant in Chinese and Japanese but only 457 (out of 14,210) samples in Arabic.
 To expand the Arabic set, we apply the same filtering process to around 2,000 samples from CIDAR Alyafeai et al.
 (2024), a human-written Arabic instruction dataset, resulting in 162 relevant samples.
-Figure 3: Overall coverage of our CARE benchmark per cultural category.
-CARE provides 16.6$\times$ more coverage of social norm and cultural commonsense questions compared to two existing instruction tuning datasets and four cultural knowledge bases (§3.1).
-[…]
-Moreover, our analysis reveals that existing resources provide highly unbalanced coverage of various cultural aspects, with a notable lack of data on social norms and commonsense, as shown in Figure 3.
+We also examine four existing cultural knowledge bases, namely, CultureBank Shi et al.
+(2024), CulturalBench Chiu et al.
+(2024b), GeoMLAMA Yin et al.
+(2022), and FORK Palta and Rudinger (2023).
+Since these datasets are exclusively in English, we manually translate them into the corresponding native language.
+We observe that these knowledge bases focus on broad regional coverage, but supply limited samples per culture: namely, 60, 61, and 50 samples for Chinese, Arab, and Japanese cultures.
+Besides, since these resources mainly use a multiple-choice or text-infilling format, we converted all answers to free-text responses, where we added further explanations while reducing stereotypes and overgeneralization.
 We manually classify all questions into one of five categories: (1) Cultural Entities, where the question asks about culture-specific entities Naous and Xu (2025), (2) Cultural Opinion, where the question asks about a subjective interpretation for a cultural entity, (3) Social Norms, where the question is about social interactions between individuals Huang and Yang (2023), (4) Cultural Commonsense, where the question is about daily phenomena that locals may take for granted Shen et al.
-(2024), and (5) Literacy, where the question is about language usage, proverbs, or slang Wuraola et al.
+(2024), and (5) Literacy, where the question is about the language, proverbs, or slang Wuraola et al.
 (2024).
-Figure 2 shows examples for these cultural aspects.
-Inter-annotator agreement is 0.90, measured by Cohen’s Kappa, for this classification.
-Finally, aligning LMs using preference optimization algorithms requires human judgments on both suitable and unsuitable responses, beyond only the existing ground-truth demonstrations.
-3.2 Multilingual Cultural Alignment Data
-To address the aforementioned shortcomings, we collect new non-English data for cultural alignment as follows:
+After filtering all resources, we found an imbalanced coverage across categories, with a notable lack of data on social norms and commonsense, as shown in Figure 2.
+To align LMs via preference optimization, we also require human judgments of both appropriate and inappropriate responses, which are not offered by existing resources.
+3.2 Multilingual Cultural Preference Data
+To address the aforementioned shortcomings, we collect new data in three languages as follows:
 Social Norm and Commonsense.
-To expand the samples on social norms and cultural commonsense, we ask native Chinese and Arabic speakers who are international college students undergoing culture transfer experiences themselves to curate such samples manually.
-To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad, culture shock, first time to the US, etc.”.
-Taking inspiration from such discussions, the speakers create 190 samples in Chinese and 196 samples in Arabic.
-This human-curated data is more authentic than synthetic data, preventing possible inaccuracy and overgeneralization (e.g.
-“In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
-Culture-specific Human Preference.
-Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment of LMs through preference optimization.
-To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs for each question within CARE.
+To expand the samples on social norms and cultural commonsense, we ask native Chinese, Arabic, and Japanese speakers who are international college students undergoing culture transfer experiences themselves to curate such samples.
+To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad”, “culture shock”, “first time to”, etc.
+Taking inspiration from such discussions, the speakers create 190, 196, and 260 samples in Chinese, Arabic, and Japanese.
+This human-curated data is more authentic than synthetic data, preventing some inaccuracy and overgeneralization (e.g., “In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
+Culture-specific Human Preference Judgments.
+Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment through preference optimization.
+To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs to each question within CARE.
 We specifically use the instruct version of recent multilingual LMs of Llama3.3-70B Dubey et al.
 (2024), Qwen2.5-72B Yang et al.
-(2024), Gemma2-27B Team et al.
+(2024a), Gemma2-27B Team et al.
 (2024), Mistral-Large, and GPT-4o.
 We also use their smaller-sized Llama3.1-8B, Qwen2.5-7B, Gemma2-9B, and Mistral-7B.
-We then instruct the annotators to rank the generated responses for each question from the most to least culturally appropriate and assign each response a rating on a scale of 1 to 10 (1: poor $\rightarrow$ 10: excellent).
-Those ratings reflect how well responses match the cultural expectations of native speakers, which we use to construct preference pairs for cultural alignment in our experiments (§4.1).
-For responses that deviate from answering the questions (e.g., generating similar questions or repetitive content), annotators do not assign a rating.
-Human Annotation.
-We hire two native Chinese speakers and two native Arabic speakers who are college students to collect and annotate all the data in CARE, as described in §3.1 and §3.2.
-Besides the 5-class cultural categories, we also annotate each sample with the Associated Culture for experiments in §4.5: Native (questions about the native culture; i.e.
-Chinese or Arab), Foreign (questions about other cultures), or General (questions that are not specific to a particular culture).
-For further fine-grained regional cultural knowledge evaluation (Appendix D.1), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide).
+We instruct annotators to rank the generated responses from the most to least culturally appropriate and assign a rating on a 1–10 scale (1: poor $\rightarrow$ 10: excellent).
+These ratings reflect how well responses align with the cultural expectations of native speakers and are used to construct preference pairs for cultural preference learning (§4.1).
+Culture $\alpha$ $r$ ${\rho}$ $\tau$
+Arab 0.86 0.90 0.88 0.71
+Chinese 0.84 0.89 0.89 0.75
+Japanese 0.92 0.93 0.92 0.77
+Table 2: Inter-annotator agreement based on Krippendorff’s $\alpha$, Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$.
+Human Annotators and Agreement.
+To curate the samples in CARE, we recruited 2 Chinese, 2 Arab, and 5 Japanese native speakers who are familiar with the respective cultures.
+These annotators manually filtered the aforementioned resources and searched online platforms for additional samples, capturing broader community knowledge and ensuring that CARE reflects cultural experiences beyond our annotator pool.
+The annotators then rated the responses to the questions in CARE.
+To ensure the quality of our preference ratings, we additionally recruited 4 Chinese, 1 Arab, and 4 Japanese annotators, who were not involved in the data curation process, to perform double annotation.
+Each additional annotator was given 75 randomly sampled questions (15 per cultural category), for which they were asked to provide their ratings of model responses.
+We then compare the ratings of the additional annotators to our initial set of annotators.
+As shown in Table 2, we find substantial inter-annotator agreements across various metrics, indicating consistent response preferences for questions within our considered cultural categories (see Appendix A.2 for per-category agreements).
+Chinese Arabic Japanese Model Entities Opinion Norms C.
+sense Literacy Average Entities Opinion Norms C.
+sense Literacy Average Entities Opinion Norms C.
+sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 4.46 6.50 5.67 6.50 2.33 5.09 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 4.40 6.60 5.80 6.67 2.93 5.28 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 2.20 3.93 2.47 2.57 1.97 2.63 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 3.20 4.67 2.80 3.20 1.67 3.11 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 2.13 5.67 4.13 4.60 2.37 3.78 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 2.77 5.73 4.30 5.17 1.90 3.97 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 1.70 3.96 2.40 2.48 1.20 2.34 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 1.83 3.90 2.43 2.33 1.23 2.35 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese, Arab culture in Arabic, and Japanese culture in Japanese.
+Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image].
+For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural preference learning (colored plot).
+Besides the 5-class cultural categories, we also label each sample with the Associated Culture for experiments in §4.5: Native (questions about the local culture; i.e., Chinese, Arab, or Japanese), Foreign (questions about other cultures; e.g., US, German, etc.), or General (questions that are not specific to a particular culture).
+For fine-grained regional cultural knowledge evaluation (Appendix D.2), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide).
 See Appendix A.1 for the annotation guideline.
 Data Split.
-For the experiments in §4 and §5, we construct culture-specific test sets with 150 questions each for Chinese and Arab cultures, randomly sampling 30 questions from each of the five cultural categories.
-The remaining 1,513 Chinese and 644 Arabic questions, along with their responses in CARE, are used to form the training set.
-This training data also includes questions about other foreign cultures beyond Chinese and Arab.
-Chinese Arabic Model Entities Opinion Norms C.
-sense Literacy Average Entities Opinion Norms C.
-sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 [Uncaptioned image] [Uncaptioned image]
-Table 1: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese and Arab culture in Arabic.
-Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image].
-For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural alignment (colored plot).
-“C.
-sense” stands for “Commonsense”.
-[…]
-Table 2: Performance comparison w.r.t. different preference learning algorithms on CARE[Uncaptioned image] data. Results show the average score improvements over the vanilla model.
+For the experiments in §4 and §5, we construct culture-specific test sets of 150 questions each for Chinese, Arab, and Japanese cultures, sampling 30 questions per cultural category.
+Remaining CARE questions (1,513 Chinese, 644 Arabic, 757 Japanese), along with rated responses, form the training set.
+This training data also includes questions about other cultures (e.g.
+US, German, etc.) beyond Chinese, Arab, and Japanese.
 4 Aligning LMs for Cultural Awareness
-Enabled by CARE, we investigate whether a small amount of high-quality cultural alignment data can improve existing models through preference optimization.
-We first conduct cultural alignment on medium-sized LMs and show how much improvement we can achieve with CARE in §4.1 and §4.2.
-Next, we examine the impact of scaling data and model size (§4.3) and whether cultural alignment sacrifices LMs’ capabilities on general tasks (§4.4).
-Finally, we analyze the importance of including data from diverse cultures during alignment (§4.5).
+We investigate whether, and to what extent, high-quality culture-specific data can enhance LMs.
+We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE that fit in 8 Nvidia A40 GPUs.
 4.1 Experiment Setup
-We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE on up to 8 Nvidia A40 GPUs.
 Human Preference Optimization.
@@
 We use the highest- and lowest-scored responses from the human preference data collected in CARE.
-We find that all three offline preference algorithms perform similarly well, as shown in Table 3.2; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.
+We find that all three offline preference algorithms perform similarly well, as shown in Table 4; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.
+[…]
+Table 4: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese, Arab, and Japanese cultures for a variety of prompting approaches, supervised fine-tuning, and preference learning using culture-specific (CARE) vs.
+general instruction-tuning (multilingual Alpaca) and preference (translated and filtered OpenOrca/UltraFeedback) data.
+SFT is performed on the instruction data only, while preference learning is conducted on the preference pairs.
 Baselines.
-First, we analyze the effectiveness of learning from cultural preference compared to general preference.
-We align LMs using existing human preference datasets, such as OpenOrca Lian et al.
-(2023) and UltraChat Bartolome et al.
-(2023), which focus on truthfulness, honesty, and helpfulness.
-For a fair comparison, we randomly sample an equal number of training pairs from these datasets as in CARE’s training set.
-Second, we examine the effectiveness of SFT on CARE versus Alpaca Cui et al.
-(2023); Chen et al.
-(2023), a general instruction-tuning dataset with Chinese and Arabic adaptations.
-Third, for prompting-based approaches, we explore CoT prompting Wei et al.
-(2022) and role-play prompting Kong et al.
+We compare cultural preference tuning on CARE with four groups of baselines: (1) General-domain preference datasets: we adopt the multilingual translations 2A2I (2024); Pan (2024); Hwang (2024) of two synthetic English preference datasets, OpenOrca Intel (2024) (2.1k Ar, 2.0k Zh, 12.9k Ja pairs) and UltraFeedback Cui et al.
+(2023a) (2.0k Ar, 1.7k Zh pairs).
+We also compare with HelpSteer3 Wang et al.
+(2025b), a human-annotated multilingual preference dataset (1.1k Zh, 0.2k Ja pairs).
+All focus on truthfulness, honesty, and helpfulness.
+(2) Instruction-tuned models: we conduct SFT on our CARE and the multilingual Alpaca corpus Cui et al.
+(2023b); Chen et al.
+(2023); fujiki (2023).
+(3) Multilingual preference tuning: MAPO She et al.
 (2024).
-See prompts used in Appendix B.
+(4) Prompting methods: CoT Wei et al.
+(2022) and role-play Kong et al.
+(2023).
+Implementation details are in Appendix B.
 LM-as-a-judge Evaluation and its Reliability.
-For evaluation, we adopt the LM-as-a-judge Zheng et al.
-(2023) strategy where we prompt a powerful LM to evaluate how aligned the generated response is to the human reference response.
-We use GPT-4o as the LM judge and prompt it to provide reasoning and a score on a scale of 1 to 10.
-We validate this setup by comparing it to the ratings of the native-speaking human annotators, where we achieve a high Pearson correlation of 0.93 between human and GPT-4o ratings.
-More details and our evaluation prompts can be found in Appendix C.
-[…]
-Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese and Arab cultures for a variety of prompting approaches, fine-tuning, and alignment using cultural (CARE) vs.
-generic instruction-tuning (Multilingual Alpaca) and preference (OpenOrca/UltraChat) data.
-SFT is performed on the instruction data only, while alignment conducts DPO on the human preference pairs.
-Aligning with cultural preferences enhances cultural awareness more effectively than using general preferences, with this improvement remaining consistent across various prompting setups.
+For evaluation, we adopt the LM-as-a-judge strategy Zheng et al.
+(2023), prompting GPT-4o to generate a rationale and assign a 1-10 score on how aligned the response is to the human reference.
+We validate this setup by comparing GPT-4o’s ratings with native annotators’, where we achieve a high Pearson correlation of 0.93.
+More details and full evaluation prompts can be found in Appendix C.
 4.2 Main Results
-Table 1 presents the results of medium-sized LMs before and after cultural alignment using CARE, while Table 3 compares these results with various baselines. We have the following key findings:
-Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) in both Chinese and Arabic compared to the vanilla checkpoints, even for these already instruction-tuned models.
+Table 3 presents the results of medium-sized LMs before and after cultural preference learning with CARE, while Table 4 compares results with various baselines. We see the following key findings:
+Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) compared to the vanilla checkpoints across all three cultures.
 This resonates with the findings of Zhou et al.
@@
 We also see a noticeable gap among LMs developed by different regions.
-The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) overall by aligning with CARE.
+The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) by conducting preference optimization with CARE.
 Interestingly, the aligned version of Gemma2-9B outperforms Qwen2.5-7B on social norms and commonsense for Chinese (more on this in §5.2).
-We also test its generalization capabilities on other culture-related tasks in Appendix D.
-Cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs.
-Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across all five cultural categories after alignment.
-However, alignment shows limited gains where the model’s initial performance is poor.
+Cultural preference learning strengthens stronger LMs but fails to address weaknesses in weaker LMs.
+Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across cultural categories after preference learning.
+However, preference learning shows limited gains where the model’s initial performance is poor.
 Mistral-7B is not improved on Chinese entities and literacy, where its starting scores are only 3.03 and 2.43.
-Similarly, Qwen2.5-7B, does not benefit from alignment on Arabic entities.
-This suggests that base models need a foundational level of cultural knowledge for alignment to be effective.
+Similarly, Qwen2.5-7B, does not benefit from preference learning on Arabic entities.
+Models also show limited progress on Japanese data, reflecting their shallow exposure to Japanese culture.
+Similarly, when the model possesses foundational cultural knowledge, utilizing the preference signals from the model itself via methods like MAPO rather than directly from human preferences can also lead to improvements.
+All these suggest that base models need basic cultural knowledge for preference learning to be effective.
 Culture-specific human preference data is helpful beyond general human preference.
-Table 3 shows that alignment with CARE’s cultural preference consistently outperforms general preference datasets by up to 42%.
-Additionally, cultural alignment with CARE brings improvement in both CoT and role-play prompting.
-While culture-specific SFT outperforms generic instruction tuning, alignment yields greater gains, highlighting the importance of human cultural preference data over ground-truth demonstrations alone.
-4.3 How does scaling affect alignment?
-We examine how scaling both model size and data volume influences the performance before and after cultural alignment.
+Table 4 shows that tuning with CARE’s cultural preference data improves performance over much larger general preference datasets, even though they contain 3–17$\times$ more training pairs.
+Cultural preference learning with CARE also improves performance in both CoT and role-play prompting setups.
+While culture-specific SFT outperforms generic SFT, preference tuning yields greater gains, highlighting the value of human cultural preferences over ground-truth demonstrations alone.
+Figure 3: Impact of model size and preference data volume on cultural awareness performance.
+The average scores (1: poor $\rightarrow$ 10: excellent) of aligned models are plotted against different % of preference pairs in CARE.
+Improvements are achieved across different model sizes and data sizes, in comparison to the vanilla model (0%).
+4.3 Scaling of Data and Model Sizes
+We examine how scaling both model size and data volume influences the performance before and after cultural preference learning.
 Specifically, we consider the Qwen2.5 {1.5B, 3B, 7B, 14B} and the Gemma2 {2B, 9B} series, which offer multiple model sizes.
@@
 Even a relatively small amount (e.g., 25%) of data in CARE can lead to improvements, particularly for smaller-sized LMs.
-Scaling up the data leads to progressively better performance, highlighting the potential benefits of employing more cultural preference data in future work.
-Figure 4: Impact of model size and preference data volume on cultural awareness performance.
-The average scores (1: poor $\rightarrow$ 10: excellent) of aligned models are plotted against different % of preference pairs in CARE.
-Improvements are achieved across different model sizes and data sizes, in comparison to the vanilla model (0%).
-4.4 Does alignment impact general capability?
-We then examine whether cultural alignment impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, and English: ArabicMMLU Koto et al.
-(2024), ChineseMMLU Li et al.
-(2023), MMLU Hendrycks et al.
-(2020), TruthfulQA Lin et al.
-(2021), and WinoGender Rudinger et al.
+Scaling up the data leads to progressively better performance, highlighting the benefits of employing more cultural preference data in future work.
+4.4 Cultural Generalization
+We evaluate CARE-aligned models on other cultural tasks using culture-related data that is not included in CARE.
+Specifically, we use the short answer questions within the Blend Myung et al.
+(2024) benchmark and multi-choice questions within the Include Romanou et al.
+(2025) benchmark, both written in the native languages.
+From Figure 4, we observe that preference learning with culture-specific human preference can also benefit out-of-domain cultural tasks, indicating the generalization capability of CARE.
+Besides, we also test the impact of cultural preference learning on general NLP benchmarks in Appendix D.3, showing that cultural preference learning does not hinder models’ overall capabilities.
+4.5 Multi-cultural Alignment
+We analyze the impact of incorporating data about different cultures in preference learning.
+To ensure a fair comparison, we select an equal number of samples for each trial.
+Specifically, we align Llama3.1-8B-Instruct with 675 Chinese and 547 Arabic pairs sampled from three different contexts: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign).
+Results are shown in Figure 5.
+Tuning on foreign-culture pairs alone lowers native performance; adding native pairs raises them (Chinese 4.69 $\rightarrow$ 4.87, Arab 2.98 $\rightarrow$ 3.50), and combining native and foreign pairs further improves performance, indicating that geographically diverse preference data can further strengthen cultural awareness.
+Figure 4: Accuracy on the Chinese, Arabic, and Japanese subsets of the short answer questions in Blend Myung et al.
+(2024) and multi-choice questions in Include Romanou et al.
 (2025).
-We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraChat and OpenOrca datasets (totaling 3,718 samples), the culturally aligned LM with cultural human preference from CARE, and the aligned LM with both general and CARE’s cultural preference.
-The results in Figure 5 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural alignment does not hinder the model’s overall capabilities.
-Figure 5: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese and Arabic) and English.
-Slight differences are observed among the LM variants after applying different alignment setups.
-4.5 Should we align with only native data?
-We analyze the impact of incorporating data about different cultures in the alignment process.
-To ensure a fair comparison, we select an equal number of samples for each trial.
-Specifically, we align Llama3.1-8B-Instruct with 676 Chinese and 546 Arabic pairs sampled from three different contexts in CARE: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign).
-Results are shown in Figure 6.
-We find that performance drops on native culture when the model is aligned using only foreign culture samples, but improves when aligned with the same native culture samples (4.69 $\rightarrow$ 4.87 in Chinese and 2.98 $\rightarrow$ 3.50 in Arabic); interestingly, performance is further improved when the model is aligned by combining native and foreign cultural data.
-This indicates that expanding geographical coverage in alignment could further enhance cultural awareness.
+CARE preference tuning improves out-of-domain cultural tasks for most cultures and models, except on the Arabic subset of Include.
+Blend does not provide a Japanese subset.
+Figure 5: Average scores of Llama3.1-8B-Instruct on local and foreign cultures when aligned using data from native, foreign, or mixed cultures.
+The highest performance on local culture is achieved when mixing local and foreign samples during preference learning.
+4.6 Language-specific LLMs
+We also explore the combined benefits of both language-specific training and cultural preference tuning.
+In particular, we conduct DPO tuning on CARE’s Japanese training split with the Swallow model variants (Llama3.1-Swallow-8B, Gemma2-Swallow-9B, and Mistral-Swallow-7B), which are continuously pre-trained and instruction-tuned on the Japanese corpus Fujii et al.
+(2024).
+Table 6 shows that language-specific continual pre-training improves each base model by 1–3 points (on a 10-point scale), while subsequent preference tuning on CARE contributing an additional 0.2–0.4 points.
+Combined, these steps enable smaller models to outperform much larger LLMs.
+For example, aligned Llama-Swallow-8B outperforms vanilla Llama-70B (6.01 vs.
+5.82), and aligned Gemma-Swallow-9B surpasses Gemma-27B (6.61 vs.
+5.85).
+The results echo our findings in §4.2: cultural alignment is consistently helpful, even when base models have enough language proficiency.
 […]
 Table 5: Average scores on CARE of larger LMs and web content, evaluated by the judge LM across all samples (1: poor $\rightarrow$ 10: excellent).
-“Human Awareness (%)” indicates the percentage of questions for which natives know the correct answer.
-Qwen2.5 developed by Chinese-based Alibaba excels on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense.
-For social norms and cultural commonsense, where native humans show high confidence, Google search often returns fewer relevant answers.
+“Human Awareness (%)” indicates the percentage of questions for which locals know the correct answer.
+Chinese-developed Deepseek-v3 and Qwen2.5 excel on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense.
+For social norms and cultural commonsense, where local humans show high confidence, Google search often returns fewer relevant answers.
+Model Entities Opinion Norms C.
+sense Literacy Avg.
+Llama family Llama3.1-8B 2.20 3.93 2.47 2.57 1.97 2.63 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 3.20 4.67 2.80 3.20 1.67 3.11 $\hookrightarrow$ Swallow-8B 5.63 6.40 5.47 6.33 4.63 5.69       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 5.77 6.57 5.60 6.70 5.40 6.01 Llama3.3-70B 6.17 6.07 4.77 6.13 5.97 5.82 $\hookrightarrow$ Swallow-70B 7.20 7.37 7.13 7.53 7.27 7.30 Gemma family Gemma2-9B 4.46 6.50 5.67 6.50 2.33 5.09 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 4.40 6.60 5.80 6.67 2.93 5.28 $\hookrightarrow$ Swallow-9B 6.07 7.27 6.90 7.20 5.10 6.51       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 6.10 7.13 7.37 6.97 5.50 6.61 Gemma2-27B 5.37 7.07 6.50 6.47 3.83 5.85 $\hookrightarrow$ Swallow-27B 6.37 7.43 6.43 7.43 6.00 6.73 Mistral family Mistral-7B 1.70 3.96 2.40 2.48 1.20 2.34 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 1.83 3.90 2.43 2.33 1.23 2.35 $\hookrightarrow$ Swallow-7B 2.70 4.43 4.07 3.87 2.00 3.41       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 3.13 4.00 4.50 4.90 2.53 3.81 Mistral-8×7B 2.10 4.37 3.23 3.30 2.30 3.06 $\hookrightarrow$ Swallow-8×7B 4.33 5.40 5.30 5.87 2.90 4.76
+Table 6: Average scores on CARE’s Japanese test set for (i) each backbone, (ii) each backbone further tuned on CARE, (iii) its Japanese-pretrained Swallow variant, and (iv) that variant further tuned on CARE.
 5 Prompting LMs for Cultural Awareness
-Figure 6: Average scores of Llama3.1-8B-Instruct on native and foreign cultures when aligned using data from native, foreign, or mixed cultures.
-The highest performance on native culture is achieved when mixing native and foreign samples during alignment.
-Beyond improving medium-sized LMs with culture-specific human preference, we evaluate the performance of larger LMs and compare with native human familiarity and web-based sources (§5.1).
-To better understand these differences, we examine the role of querying language (§5.2).
+We evaluate larger LMs and compare to human familiarity and web-based sources (§5.1), as well as querying in different languages (§5.2).
 5.1 Main Results
@@
 Table 5 shows the results.
-Larger models are better for both Chinese and Arabic, with GPT-4o achieving the best overall performance.
-The more Chinese-centric model Qwen2.5-72B excels in the knowledge about cultural entities (8.42), social norms (9.06), and literacy (8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture.
+Larger models are better for all three cultures, with closed-source ones like GPT-4o and Deepseek-v3 leading the overall performance.
+The more Chinese-centric model Deepseek-v3 and Qwen2.5-72B excels in the knowledge about cultural entities (8.96 and 8.42), social norms (9.03 and 9.06), and literacy (9.03 and 8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture.
 This is interesting and likely because cultural commonsense knowledge is often unstated in the native language, while more balanced and thoughtful opinions may be expressed in foreign languages.
-To illustrate and test this hypothesis, we conduct further analysis in §5.2.
-Human Awareness.
-To assess human familiarity with the culture-specific questions, we ask native-speaking annotators whether they knew the correct answers without looking them up when presented with questions from CARE.
-Table 4 shows the percentages of questions humans could answer immediately with confidence, which are very high for social norms and commonsense knowledge.
-Search Engine.
-To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google [^1] Google programmable search engine API: https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE.
-Results are shown in Table 4.
-Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.
+We further test this hypothesis in §5.2.
 (a) Average scores achieved by Qwen2.5-72B-Instruct and GPT-4o when prompted in native languages versus English.
@@
 However, the gaps narrow or sometimes reverse for cultural norms and commonsense.
+Human Awareness.
+To assess human familiarity with the culture-specific questions, we ask native-speaking annotators whether they knew the correct answers without looking them up when presented with questions from CARE.
+Table 5 shows the percentages of questions humans could answer immediately with confidence, which are very high for social norms and commonsense knowledge.
+Search Engine.
+To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google111 https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE.
+Results are shown in Table 5.
+Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.
 5.2 Cross-lingual Analysis
-We translate all questions and responses from CARE that are written in Chinese and Arabic into English to enable cross-lingual evaluation.
-Figure 7(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than for Arabic.
+We translate all questions and answers in CARE that are written in Chinese, Arabic, and Japanese into English to enable cross-lingual evaluation.
+Figure 6(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than Arabic and Japanese.
 This advantage is also especially obvious for the literacy category, while the gap narrows in categories such as cultural commonsense.
-A possible explanation is that English corpora contain fewer discussions about Chinese or Arab cultures.
 We also examine how language choice impacts search engine results.
@@
 Interestingly, this trend reverses for questions related to cultural commonsense and social norms, such as “In China, does the ticket time indicate when the feature movie starts, or are there trailers played before the main feature?”, where English searches yield higher-quality content.
-This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, whereas non-native speakers are more likely to actively seek out and document it.
-A similar pattern is observed for Arabic.
-For a small number of Arabic questions (10% opinion, 14% social norms, 14% commonsense, and 7% literacy), the search engine also failed to return any results.
-One such example is “A Jordanian friend mentioned that it’s polite not to open a gift in front of the giver.
-Is this true?”.
+This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, but non-native speakers are more likely to seek out information or document it.
 6 Conclusion
-We introduce CARE, a human-written, multilingual resource comprising 2,580 culture-specific questions about Chinese and Arab cultures, along with responses that include human preferences.
-Through extensive experiments, we investigate that a small amount of high-quality cultural alignment data can improve existing LMs via preference optimization, across different model families and scales.
-This cultural-specific alignment offers benefits that go beyond generic human preference and proves effective under various prompting setups.
-Our analyses also reveal gaps in cultural awareness among LMs, native-speaking humans, and search engines.
-For example, cultural commonsense is often left implicit in the native language but is more explicitly mentioned in foreign-language content.
-By releasing CARE, we hope to encourage the development of more inclusive and culturally adaptive LMs.
+We introduce CARE, a multilingual, human-written resource comprising 3,490 culture-specific questions about Chinese, Arab, and Japanese cultures, along with human-rated responses.
+Through extensive experiments, we investigate that a small amount of high-quality cultural preferences can improve LMs via preference optimization, across various model families and scales.
+Our analyses reveal gaps in cultural awareness among LMs, native speakers, and search engines, especially regarding implicit cultural commonsense.
+By releasing CARE, we hope to foster the development of more culturally adaptive LMs.
 Limitations
-In this work, as we investigate how to improve LMs by aligning with multilingual multicultural human preference data, we also identified a couple of areas for further exploration.
-First, while filtering the Arabic portion of the Aya dataset, we notice a large number (3,831) of questions centered on Islamic knowledge and religious history.
-We exclude them, as religion-specific understanding in LMs falls outside the scope of our study that focuses on culture-specific phenomena and factual knowledge, as defined in §3.1.
-Nonetheless, future research on religious QA or alignment could adopt a similar approach to CARE by collecting preference data from credible religious sources.
-Second, LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding.
+During data collection, we intentionally excluded highly sensitive and controversial topics (e.g., “Are security concerns about the niqab exaggerated or justified?”), as these tend to elicit divergent personal views.
+Our work addresses generally shared topics within each cultural group (e.g., “Do Chinese people like stinky tofu?” in the Opinion category of CARE), allowing us to capture collective preferences that reflect broader cultural consensus Bakker et al.
+(2022) with a neutral point of view (e.g., “Some Chinese like stinky tofu.
+It is a traditional food made by $\ldots$ It is most popular in $\ldots$ regions.
+At the same time, many Chinese people don’t like it $\ldots$” ).
+This approach facilitates more objective and replicable evaluation, as evidenced by the high inter-annotator agreement reported in §3.2, thus supports more reliable algorithmic comparisons, which are the primary focus of this paper.
+The scope of our study is complementary to other work  Chiu et al.
+(2024a); Kirk et al.
+(2024); Huang et al.
+(2025) that explored diverse values and subjective viewpoints Zhong et al.
+(2021).
+{CJK*}
+UTF8gbsn We also note that LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding.
 For instance, when asked “In China, what does the term “laowai” mean when referring to foreigners?”, LMs often incorrectly interpret the term as disrespectful.
-In reality, “laowai” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations.
-While CARE includes examples that clarify such misunderstandings through human-annotated responses, future studies in cultural red-teaming can further investigate and address these failure modes.
+In reality, “laowai (老外)” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations.
+While CARE includes examples that clarify such misunderstandings through human-written responses, future studies in cultural red-teaming can further investigate and address these failure cases.
+Ethics Statement
+In this study, we employed in-house annotators to collect cultural samples and provide preference judgments in CARE.
+The annotators for the Chinese and Arabic samples were university-level students fluent in the respective languages.
+Japanese annotators were Japanese workers with university degrees and cross-cultural experience abroad.
+Each Chinese and Arabic annotator was paid at \$18 per hour, exceeding the U.S.
+federal minimum wage.
+We ensured that no personal data was collected from the annotators and emphasized that participation was voluntary.
+We also informed the annotators that the collected data would be used to enhance the cultural awareness of LMs.
+The culture-related questions and reference responses in CARE are either from the existing works or human-written sentences provided by our annotators.
 Acknowledgments
-The authors would like to thank Nour Allah El Senary, Jad Matthew Bardawil, Chao Jiang, and Yuming Pan for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback.
+The authors would like to thank Chao Jiang, Jad Matthew Bardawil, Nour Allah El Senary, Ruohao Guo, Xiaofeng Wu, Yuming Pan, and Zirui Shao for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback.
 We would also like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work.
@@
 A.1 Annotation Details
-To construct the question-answer pairs in CARE, we recruited two native Chinese speakers and two native Arabic speakers who are undergraduate students.
-The annotators were paid at the rate of 16 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs.
-The detailed annotation guideline for dataset construction is provided in Figures 16 to 20, and the instruction for preference annotation is provided in Figures 14.
+To construct the question-answer pairs in CARE, we recruited native Chinese, Arabic, and Japanese speakers as mentioned in § 3.2.
+Among our annotators, 5 Japanese annotators were workers with experience living abroad (e.g., in the US), while the rest were international college students.
+Annotators were paid at the rate of 18 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs.
+The detailed annotation guideline for dataset construction is provided in Figures 15 to 19, and the instruction for preference annotation is provided in Figures 13.
 We also show our rank-and-rate web interface for collecting human cultural preference pairs in Figure 14.
-A.2 Statistics
-Table 5 shows the detailed statistics for Chinese and Arab culture in CARE across the cultural categories, stratified by the data source.
-The questions specific to Chinese culture are written in Chinese, while the questions specific to Arab culture are written in Arabic.
+A.2 Annotation Agreements
+To assess annotators’ agreement on this culture preference judgment task, we invite additional native speakers for each culture, and report the Pearson correlation on different cultural categories in Table 11.
+It shows that high correlation is achieved among annotators across five cultural categories.
+Culture Entities Opinion Norms Commonsense Literacy
+Arab 0.96 0.90 0.84 0.88 0.97
+Chinese 0.84 0.90 0.96 0.94 0.99
+Japanese 0.99 0.88 0.96 0.85 0.99
+Table 11: Inter-annotator agreement measured by Spearman’s $\rho$ on five cultural categories. Agreement for this preference judgment task is consistently high across different cultural categories.
+A.3 Statistics
+Table 8 shows the detailed statistics for Chinese, Arab, and Japanese cultures in CARE across the cultural categories, stratified by the data source.
+The questions specific to Chinese, Arab, and Japanese cultures are written in Chinese, Arabic, and Japanese, respectively.
 The culturally-relevant samples obtained from the instruction-tuning dataset mostly fall within the cultural entities, cultural opinion, or literacy categories, while samples obtained from cultural knowledge bases provide more social norms and cultural commonsense data.
@@
 These samples were obtained when filtering the instruction tuning datasets and cultural knowledge bases (§3.1).
-We wrote these foreign samples in both Chinese and Arabic versions and used them as part of our training set and in our analysis on the impact of the source culture in alignment (§4.5).
+We wrote these foreign samples in the native language and used them as part of our training set and in our analysis on the impact of the source culture in preference learning (§4.5).
 We also manually collected more samples for social norms and cultural commonsense from online websites to cross 100 samples in each category.
-We ensured these foreign samples do not include anything relevant to Chinese or Arab cultures.
-Table 7 shows the detailed statistics for Chinese and Arabic samples in CARE across the geographic scope.
+We ensured these foreign samples do not include anything relevant to Chinese, Arab, or Japanese cultures.
+Table 10 shows the detailed statistics for Chinese, Arabic, and Japanese samples in CARE across the geographic scope.
 Most samples fall into nationwide category, while sub-nationwide samples provide insights about more detailed cultural information, continent-wide and worldwide samples assess understanding between several cultures.
+A.4 Comparison with existing datasets
+Table [Uncaptioned image]CARE: Assessing the Impact of Multilingual Human Preference Learning on Cultural Awareness presents illustrative examples from representative datasets that aim to improve LMs’ cultural awareness.
+While CultureLLM Li et al.
+(2024a) focuses on synthetic data without preference supervision, and Prism Kirk et al.
+(2024) provides human ratings on generic English questions, neither provides proper multilingual and culture-specific contexts with human preferences.
+In contrast, CARE provides human preferences on culturally grounded topics across multiple languages.
 Appendix B Implementation Details
-DPO Alignment.
-We perform preference optimization using DPO for 3 epochs until the loss converges.
+SFT Tuning.
+We conduct full-parameter SFT on the instruction data only, using the questions and human-written reference responses in CARE.
+We tune the learning rate in the range $\{1e^{-5},2e^{-5}\}$, and train with a batch size of 128 on 4–8 NVIDIA A40 GPUs.
+Preference Tuning.
+We perform full-parameter preference optimization individually for each language using DPO, KTO, and SimPO for 3 epochs until the loss converges.
 Training is done with a batch size of 128 on 4-8 NVIDIA A40 GPUs.
-We tune the learning rate in the range $\{1e^{-7},5e^{-7},9e^{-7}\}$ and set beta for DPO as 0.1.
+We tune the learning rate in the range $\{3e^{-7},5e^{-7},7e^{-7}\}$ and set beta as 0.1.
 The training involves full fine-tuning with 5 warmup steps and employs a linear learning rate scheduler.
+MAPO Tuning.
+We use the same base LLM to be tuned as the rollout model and use NLLB-600M-distilled Costa-Jussà et al.
+(2022) to calculate the alignment score as suggested in the original paper She et al.
+(2024).
+We conduct DPO tuning on the generated preference pairs and report the results of the first iteration.
 LM Inference.
-For open-sourced LMs, we run inference on 1 NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al.
+For open-sourced LMs, we run inference on one NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al.
 (2023).
@@
 For the closed-source GPT-4o LM, we run inference with Azure OpenAI API.
+Baseline preference datasets.
+Table 12 lists the URLs of the multilingual adaptations of the general-domain preference datasets used in our experiments (§4.2).
+HelpSteer3 is natively multilingual, with prompts and responses originally written in various languages and manually annotated preference labels.
+In contrast, the multilingual versions of UltraFeedback and OpenOrca were generated by automatically translating the English response pairs into the target languages, followed by automatic quality control to remove poorly translated samples.
+Their preference signals are also synthetic.
+UltraFeedback relies on GPT-4 to rank the responses, while OpenOrca assumes GPT-4’s responses are always superior to Llama’s when constructing preference pairs  Wang et al.
+(2025a).
+Dataset Language Source HuggingFace Repository
+OpenOrca ar translation 2A2I/argilla-dpo-mix-7k-arabic
+ja translation yongtae-jp/orca_dpo_pairs_ja
+zh translation wenbopan/Chinese-dpo-pairs
+UltraFeedback ar translation 2A2I/argilla-dpo-mix-7k-arabic
+zh translation wenbopan/Chinese-dpo-pairs
+HelpSteer3 zh, ja native nvidia/HelpSteer3
+Table 12: Multilingual adaptations of baseline preference datasets. “Source” indicates whether the non-English data are natively collected or derived via automatic translation.
 Baselines Prompt Templates.
@@
 Role-play Prompt Template
-You are a native [Chinese/Arab] person, familiar with
-[Chinese/Arab] culture and traditions.
+You are a native [Chinese/Arab/Japanese] person, familiar
+with [Chinese/Arab/Japanese] culture and traditions.
 {Question}
 Figure 7: Prompt template for role-play inference.
-The LM is told to take on the persona of a native [Chinese/Arab] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
+The LM is told to take on the persona of a native [Chinese/Arab/Japanese] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
 CoT Prompt Template
@@
 Figure 8: Prompt template for CoT inference. We provide the LM with the test question in {Question}, then ask it to think step by step when providing the answer.
-Chinese Arabic Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92
-Table 8: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both native culture and foreign culture.
+Chinese Arabic Japanese Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 2.50 5.89 7.20 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 6.75 5.77 6.40 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 4.25 6.57 8.00 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 4.50 6.45 6.40 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92 7.00 8.32 8.00
+Table 13: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both local culture and foreign culture.
 Appendix C LM-as-a-Judge
@@
 Figure 9: Rating score distributions of the judge LM (GPT-4o) and native human evaluators. The judge LM highly correlates with human ratings.
+[…]
+Table 14: Performance comparison w.r.t. different preference learning algorithms on CARE[Uncaptioned image] data. Results show the average score improvements over the vanilla model.
 Appendix D Additional Results
-D.1 Performance by Geographic Scope
-Table 8 shows the results achieved by different LMs when stratified by geographic scope.
+D.1 Performance by Preference Learning strategies
+Table C.2 shows per-category performance when tuned with different preference learning strategies.
+D.2 Performance by Geographic Scope
+Table 13 shows results from different LMs when stratified by geographic scope.
 We find that the models struggle the most with sub-nationwide questions.
 Most models achieve the best scores on continent-wide or worldwide questions.
-D.2 Performance on other cultural tasks
-We also evaluate the performance of LMs after cultural alignment with CARE on other cultural tasks using out-of-domain culture-related data.
-Specifically, we use the Include benchmark Romanou et al.
-(2025), as it offers multilingual versions for each culture and employs a multi-choice format, enabling us to assess the robustness of our alignment.
-From Table 9, we observe that alignment with culture-specific human preference can also benefit the performance on the Include benchmark, indicating the generalization capability of CARE.
-[Uncaptioned image] Gemma2-9B [Uncaptioned image] Llama3.1-8B [Uncaptioned image] Qwen2.5-7B [Uncaptioned image] Mistral-7B Vanilla CARE Vanilla CARE Vanilla CARE Vanilla CARE Include (zh) 57.79 59.44 51.00 54.68 75.22 75.59 40.36 40.73 Include (ar) 57.42 54.71 50.54 51.08 56.15 55.97 35.32 36.95
-Table 9: Accuracy (%) on the Include benchmark for Chinese and Arabic subsets, before and after cultural alignment with CARE.
-Improvements in the Include benchmark can also be observed after alignment with CARE.
+D.3 Performance on general NLP tasks
+We then examine whether cultural preference learning impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, Japanese, and English: ArabicMMLU Koto et al.
+(2024), ChineseMMLU Li et al.
+(2023), the Japanese subset of MMLU-ProX Xuan et al.
+(2025), MMLU Hendrycks et al.
+(2020), TruthfulQA Lin et al.
+(2021), and WinoGender Rudinger et al.
+(2018).
+We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraFeedback and OpenOrca datasets, and the culturally aligned LM with cultural human preference from CARE.
+The results in Table 15 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural preference learning does not hinder the model’s overall capabilities.
+Arabic Chinese Japanese Model Vanilla DPO (w.
+Orca+Ultra) DPO (w.
+CARE) Vanilla DPO (w.
+Orca+Ultra) DPO (w.
+CARE) Vanilla DPO (w.
+Orca) DPO (w.
+CARE) MMLU (ar/zh/ja) 56.4 57.1 56.5 55.5 55.5 55.2 28.8 28.4 30.2 MMLU (en) 69.1 68.8 69.1 69.1 68.9 69.1 69.1 68.8 69.0 TruthfulQA 37.8 38.7 38.6 37.8 38.8 40.8 37.8 41.1 38.7 WinoGender 52.6 55.0 52.8 52.6 54.4 52.1 52.6 54.7 52.9
+Table 15: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese, Arabic, and Japanese) and English.
+Slight differences are observed among the LM variants after applying different preference learning methods.
 Appendix E Responsible NLP Checklist
@@
 We acknowledge the use of GPT-4 for grammar checking and word polishing.
+Our annotators were undergraduate student employees paid at their normal hourly rate of 18 USD per hour.
+All annotators were informed that they were assessing data to be used in a research project focused on “studying the cultural awareness of various language models”.
 Cultural entities and opinion evaluation